Present and future value tables of $1 at 3% are presented below. 103000 097087 10000 097087 10300 000 2 106090 | 094260 | 2.0300 109273 0.91514 3.0909 2.82861 1.12551 0.88849 4.1836 1.15927 0.86261 5.3091 119405 0.83748 6.4684 122987 0.81309 76625 6.23028 78923 1.26677 0.78941 8.8923 701969 91347 2.09091.97087 2.91347 3.71710 4.3091 3.82861 4.71710 5.41719 6.6625 5.57971 6,41719 91591 723028 130477 076642 101591 778611 10.4639 801969 1.34392 074409 11.4639 8.53020 11807887861 138423 0.72242 12.8078 925262 131920 9.53020 3.1836 4 5 6 4.57971 15.4684 9 10 12 1.42576- -0.70138- 141920- 995400- 146178 - 10.25262- 13 1.46853 068095 5617810 63496 6.0863 10.95400 151259 0.66112 170863 1129607 17.5989 1163496 155797 | 064186 | 185989 ーー11.93794-1-191569- 12.29607 6 -6471-1 0.62317 t 201569 12.56110 207616-112.93794 15 Bill wants to give Maria a $620,000 gift in 2 years. If money is worth 6% compounded semiannually, what is Marias gift worth today?
Answer #1

We use the formula:
A=P(1+r/200)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

620000=P(1+0.06/2)^(2*2)

P=620,000/(1+0.06/2)^(2*2)

=$620,000*0.88849

=$550863.80(Approx).